# Submodular Optimization for Keyframe Selection & Usage in SLAM

### v1

@@ I Introduction
LiDAR is an increasingly popular sensing modality for Simultaneous Localization And Mapping (SLAM) because of its superior range measurement accuracy across different environments and conditions. LiDAR odometry estimation algorithms work by computing a rigid body transform that best aligns the most recent scan, i.e., point cloud, with previously registered points clouds. Given the density of LiDAR data, having a SLAM system that saves all incoming point clouds would be infeasible from both a memory and computation point of view. A more pragmatic approach is to save unique point clouds called keyframes that can be combined to form descriptive submaps for future scan alignment. Despite keyframes and submaps playing a central role in SLAM, their treatment has largely been left to heuristics which lack generality and require extensive, environment-specific tuning to achieve adequate performance. Another shortcoming of many LiDAR SLAM pipelines is their inability to create maps on-the-fly tailored for sharing with other agents or downstream processes that require specific information. We address the issues of online keyframe selection and usage by leveraging the unique properties of submodular optimization to i) save unique keyframes according to a neural network which encodes point cloud similarity and ii) generate submaps via an optimization of the keyframes which best constrain scan alignment. We also propose a principled method for generating size-constrained summary maps for sharing or downstream processes.
@@
Our key insight is that decisions regarding keyframe selection and usage in addition to map summarization can be formulated as combinatorial optimization problems. Submodularity, or the mathematically defined notion of diminishing returns, captures the trade-off inherent in these problems and provides efficient algorithms for approximately solving NP-hard combinatorial problems with suboptimallity guarantees. Submodular optimization has been used in a variety of robot applications, in particular single- and multi-agent coverage [1, 2, 3], but has been used sparingly in SLAM. The most relevant to our work is that of [4] which used submodular optimization for selecting the most useful set of features to track over a window for visual-inertial odometry. Other SLAM works that utilize submodular optimization have primarily focused on sparsifying pose graphs [5, 6].
@@ II Related Works
The decision to save a point cloud as a keyframe—a process refered to as keyframe selection—must balance localization accuracy with memory usage since keyframes are used as target point clouds for registration but must be saved indefinitely in memory [8]. LiDAR keyframe selection has largely relied on heuristics such as distance from other keyframes [9, 10], which can lead to keyframe sets that are redundant or insufficient for localization depending on the threshold parameter. A few works have selected sparse keyframe sets via feedback from scan alignment [11, 12], we use this strategy to supplement a proactive policy which selects keyframes with distinct descriptors.
Submaps are composed of nearby (local) keyframes and are important for keeping the computation for point cloud registration manageable. Despite their importance, few have investigated submap generation and its effect on LiDAR SLAM. Instead, most LIO algorithms rely on an excess of nearby keyframes [9, 10]. Past work has demonstrated the value of including keyframes from the edge of the scan [13, 14]; however, larger submaps demand more computational resources. Aside from keyframe-based methods, other direct LIO algorithms use a sliding window local map with all recent points [15, 16], but these methods include all points regardless of their relevance which inflates the submap size.
One technique that has shown to be effective at (passively) characterizing the quality of submaps/local features is the eigenvalues of the Hessian of the scan alignment optimization [17, 18, 19, 20]. However these methods only use the Hessian eigenvalues as a passive metric for monitoring localization performance. A related work used the Hessian eigenvalues to select better correspondences for Iterative Closest Point (ICP) algorithms [21] in generic point cloud alignment. A similar approach for Visual Inertial Odometry (VIO) [4] showed that optimizing a feature set over the minimum eigenvalue of an information matrix using submodular optimization yielded a nearly optimal set with low computation times.
In addition to feature selection in VIO, submodular optimization has been used in SLAM to prune LiDAR keyframes [8], generate sparse pose graphs [5, 6], and select navigation anchor points [22]. Streaming submodular algorithms [23, 24] have been used to summarize large image datasets, but have not been used for LiDAR scan summarization or SLAM to the best of our knowledge.
@@ III Preliminaries
The marginal value of a set function given an added element is defined as $\Delta f(A\cup\{e\})=f(A\cup\{e\})-f(A)$ where the element $e\in E$ is added to set $A$.
@@
A set function $f:{S}\rightarrow\mathbb{R}$ is submodular if for $A\subseteq B\subseteq E$ and any element $e\in E\,\backslash\,B$ then
$f(A\cup\{s\})-f(A)\geq f(B\cup\{s\})-f(B).$ (2)
Put simply, submodulary is equivalent to diminishing returns.
@@
We conclude with a discussion on streaming submodular algorithms which have proven to be effective at solving combintorial optimization problems with a single pass over the data rather than multiple passes with greedy. The streaming algorithm most pertinent to this work is described in Section V-C and given by Algorithm 3, but the intuition is that elements with high marginal value are added to a potential solution based on a guess of the optimal function value. With a range of optimal value guesses, the streaming algorithm can build several potential solution sets and return the best which is guaranteed to be within a 1/2-$\epsilon$ factor of the optimal set where $\epsilon$ is a granularity constant [23].
Notation. We call the ordered set of all point clouds collected during a trajectory E. We denote the ordered set of chosen keyframes as $\mathcal{K}\subseteq\textbf{E}$, and we denote the local submap as $\mathcal{S}\subseteq\mathcal{K}$. We use an upper case letter to indicate a member of a set and use the most specific set whenever possible, for example, a point cloud which is an element of the submap is also a keyframe but denoted as $S\in\mathcal{S}$. When needed, we denote the position of an element in its set using a subscript index $S_{i}$, and we adopt the notation $\mathcal{S}_{<i}$ to denote the subset of $\mathcal{S}$ such that only elements with index less than $i$ are included, e.g., $\mathcal{S}=\{S_{0},S_{1},S_{2}\},~{}\mathcal{S}_{<2}=\{S_{0},S_{1}\}$.
@@ IV LiDAR Scan Neural Network
We adopt the neural network proposed by [7]; the overall framework is presented in Fig. 2. Compared to point- and discretized-based methods [27, 28, 29], the neural network is compact and can achieve fast execution without any semantic information making it generalizable. Instead of using the overlap between the projected range images to supervise the training, we measure the similarity between two point clouds $E_{i}$ and $E_{j}$ directly using the 3D Jaccard index [11]. Defined as $J(E_{i},E_{j})={|E_{i}\bigcap E_{j}|}/{|E_{i}\bigcup E_{j}|}$, the Jaccard index is the equivalent of overlap in the 3D point cloud domain to provide a direct measure of point cloud similarity which enhances training results. Unless otherwise specified, the dimension of the output descriptor is $\text{dim}(\mathcal{G})={256}$.
@@ V-A Keyframe Selection via Feature Descriptor
Our approach relies on the neural network discussed in Section IV to efficiently determine the similarity between any two point clouds. We define the marginal gain of any point cloud as the minimum Euclidean distance in feature space between that point cloud $E$ and any of the previously selected keyframes $\Delta f(\mathcal{K}\cup E)=\min_{K\in\mathcal{K}}||\phi(E)-\phi(K)||$. This marginal value inherently defines an objective function
$\max_{\mathcal{K}\subseteq\textbf{E}}f(\mathcal{K})=\sum_{K_{i}\in\mathcal{K}}\min_{K_{j}\in\mathcal{K}_{<i}}||\phi(K_{i})-\phi(K_{j})||.$ (6)
Although we do not directly optimize Equation 6, we still obtain a suboptimality bound by only selecting keyframes with high marginal value because $f$ is a non-decreasing monotonic submodular set function, as proven in the following proposition.
@@
Algorithm 1 describes our keyframe selection strategy, where a keyframe is selected according to a feature distance condition with threshold $\alpha$, and a degeneracy condition with threshold $\beta$. In addition to a boolean variable representing whether to save a keyframe or not, we also return the added suboptimality (with respect to Equation 6) $\gamma$ such that the marginal value of the new keyframe is $\alpha-\gamma$. In line 1 we find the smallest distance to any keyframe in feature space. If the closest feature distance is greater than $\alpha$, we return true and report no additional suboptimality on line 4. On lines 5-8, we compute degeneracy and save a keyframe if it exceeds the threshold $\beta$. If neither condition is met, Algorithm 1 returns false. By knowing the marginal value of each keyframe at the time of selection, we also maintain a known suboptimality bound which is given in the following proposition.
@@ V-B Submap Generation via Submodular Maximization
$\min_{\textbf{t},\,\textbf{q}}\sum_{c\in\mathcal{C}}\mathcal{E}(\textbf{t},\textbf{q},E,T),$ (7)
@@
Assumption 1 allows us to approximate $\mathcal{H}_{tt}$ as a sum of sub-Hessians. We denote the sub-Hessian associated with keyframe $K_{i}$ as $\partial\mathcal{H}_{tt,K_{i}}$ and approximate it as $\partial\mathcal{H}_{tt,K_{i}}=\mathcal{H}_{tt}(E-\mathcal{S}_{<i},K_{i})$, where $\mathcal{S}_{<i}$ is the submap including all keyframes with index less than $i$, and the subtraction of point clouds indicates removing the intersection of points ($A-B=A-(A\cap B))$. Because each Hessian is the outer product of the Jacobian with itself, the Hessian must be positive semi-definite $\mathcal{H}_{tt}=J^{T}_{t}J_{t}\succeq 0$, which intuitively means each keyframe cannot make Equation 7 less constrained. We define any keyframe which introduces enough new correspondences such that $\partial\mathcal{H}_{tt,Ki}\succ 0$ as a constructive keyframe.
@@
$\Delta f(\mathcal{S}\cup K)=\lambda_{min}(\mathcal{H}_{tt}(E,\mathcal{S}\cup K))-\lambda_{min}(\mathcal{H}_{tt}(E,\mathcal{S}))$
$=\lambda_{min}(\mathcal{H}_{tt}(E,\mathcal{S})+\partial\mathcal{H}_{tt,K})-\lambda_{min}(\mathcal{H}_{tt}(E,\mathcal{S}))$
$=\min_{||\textbf{u}||=1}\textbf{u}^{T}[\mathcal{H}_{tt}(E,\mathcal{S})+\partial\mathcal{H}_{tt,K}]\textbf{u}-\min_{||\textbf{v}||=1}\textbf{v}^{T}\mathcal{H}_{tt}(E,\mathcal{S})\textbf{v}$
$\geq\bar{\textbf{u}}^{T}[\mathcal{H}_{tt}(E,\mathcal{S})+\partial\mathcal{H}_{tt,K}]\bar{\textbf{u}}-\bar{\textbf{u}}^{T}\mathcal{H}_{tt}(E,\mathcal{S})\bar{\textbf{u}}$
$=\bar{\textbf{u}}^{T}\partial\mathcal{H}_{tt,K}\bar{\textbf{u}}>0,$
where $\bar{\textbf{u}}$ is the minimizer of $\min_{||\textbf{u}||=1}\textbf{u}^{T}[\mathcal{H}_{tt}(E,\mathcal{S})+\partial\mathcal{H}_{tt,K}]\textbf{u}$, and the final line is true because the keyframe is constructive. The numerator of Definition 4 is the minimum marginal gain attainable, so if the function is monotone increasing, the submodularity ratio must be bounded away from 00 proving Equation 8 is approximately submodular. ∎
Algorithm 2 shows our approach to submap generation, which begins by finding and saving maximum Hessian eigenvalues (lines 3-6). Each greedy iteration (lines 9-23) computes the highest marginal value keyframe, then adds it to the submap and removes its points from the scan such that the following iterations will compute sub-Hessians. The resulting submap enjoys a suboptimality bound from Definition 5 if it reaches the cardinality constraint or achieves the maximum attainable value of Equation 8 if the marginal value reaches zero. We only approximate Hessians using a percent ($\sim 25\%$) of randomized scan points as previous work has shown this has a minimal effect on the relative magnitude of ICP Hessian eigenvalues [21].
@@
where $A$ and $B$ are Hermitian matrices with ordered eigenvalues such that $\lambda_{1}\geq...\geq\lambda_{n}$. We use Weyl’s $1st$ inequality ($i=n,j=1$) to upper bound the minimum eigenvalue of the Hessian $\lambda_{min}(\mathcal{H}_{tt}+\partial\mathcal{H}_{tt,K})\leq\lambda_{min}(\mathcal{H}_{tt})+\lambda_{max}(\partial\mathcal{H}_{tt,K})$. Next, we note that the maximum eigenvalues of the Hessians obtained on line 4 upper bound the sub-Hessian eigenvalues $\lambda_{max}(\mathcal{H}_{tt}(E,K))\geq\lambda_{max}(\partial\mathcal{H}_{tt,K})~{}\forall~{}\mathcal{S}$. This is because the maximum eigenvalue is monotone non-decreasing (Weyl’s $2nd$ inequality $i=1,j=n$) and the Hessian obtained on line 4 is the sum of the sub-Hessian with it’s counterpart formed by points shared by the scan, keyframe, and previous submap. This combined with the minimum eigenvalue upper bound imply an upper bound on the marginal value of adding keyframe $K$ of $\Delta\lambda_{min}(\mathcal{S}\cup K)\leq\,\lambda_{max}(\mathcal{H}_{tt}(E,K))$. This means any keyframe with an upper bound smaller than the best marginal gain in the current iteration can be ignored.

### v2

@@ I Introduction
LiDAR is an increasingly popular sensing modality for Simultaneous Localization And Mapping (SLAM) because of its superior range measurement accuracy across different environments and conditions. LiDAR odometry estimation algorithms work by computing a rigid body transform that best aligns the most recent scan, i.e., point cloud, with previously registered point clouds. Given the density of LiDAR data, having a SLAM system that saves all incoming point clouds would be infeasible from both a memory and computation point of view. A more pragmatic approach is to save unique point clouds called keyframes that can be combined to form descriptive submaps for future scan alignment. Despite keyframes and submaps playing a central role in SLAM, their treatment has largely been left to heuristics which lack generality and require extensive, environment-specific tuning to achieve adequate performance. Another shortcoming of many LiDAR SLAM pipelines is their inability to create maps on-the-fly tailored for sharing with other agents or downstream processes that require specific information. We address the issues of online keyframe selection and usage by leveraging the unique properties of submodular optimization to i) save unique keyframes according to a neural network which encodes point cloud similarity and ii) generate submaps via an optimization of the keyframes which best constrain scan alignment. We also propose a principled method for generating size-constrained summary maps for sharing or downstream processes.
@@
Our key insight is that decisions regarding keyframe selection and usage in addition to map summarization can be formulated as combinatorial optimization problems. Submodularity, or the mathematically defined notion of diminishing returns, captures the trade-offs inherent in these problems and provides efficient algorithms for approximately solving NP-hard combinatorial problems with suboptimality guarantees. Submodular optimization has been used in a variety of robot applications such as single- and multi-agent coverage [1, 2, 3], but has been used sparingly in SLAM. The most relevant to our work is that of [4] which used submodular optimization for selecting the most useful set of features to track over a window for visual-inertial odometry. Other SLAM works that utilize submodular optimization have primarily focused on sparsifying pose graphs [5, 6].
@@ II Related Works
The decision to save a point cloud as a keyframe—a process refered to as keyframe selection—must balance localization accuracy with memory usage. This is because keyframes are used as target point clouds for registration but must be saved indefinitely in memory [8]. LiDAR keyframe selection has largely relied on heuristics such as distance from other keyframes [9, 10], which can lead to keyframe sets that are redundant or insufficient for localization depending on the threshold parameter. A few works have selected sparse keyframe sets via feedback from scan alignment [11, 12]. We use this strategy to supplement a proactive policy which selects keyframes with distinct descriptors.
Submaps are composed of nearby (local) keyframes and are important for keeping the computation for point cloud registration manageable. Despite their importance, few have investigated submap generation and its effect on LiDAR SLAM. Instead, most LIO algorithms rely on an excess of nearby keyframes [9, 10]. Past work has demonstrated the value of including keyframes from the edge of the scan [13, 14]; however, larger submaps demand more computational resources. Aside from keyframe-based methods, other direct LIO algorithms use a sliding window local map with all recent points [15, 16]. These methods include all points regardless of their relevance, which inflates the submap size.
One technique that has shown to be effective at (passively) characterizing the quality of submaps/local features is the eigenvalues of the Hessian of the scan alignment optimization [17, 18, 19, 20]. However, these methods only use the Hessian eigenvalues as a passive metric for monitoring localization performance. A related work used the Hessian eigenvalues to select better correspondences for Iterative Closest Point (ICP) algorithms [21] in generic point cloud alignment. A similar approach for Visual Inertial Odometry (VIO) [4] showed that optimizing a feature set over the minimum eigenvalue of an information matrix using submodular optimization yielded a near optimal set with low computation times.
In addition to feature selection in VIO, submodular optimization has been used in SLAM to prune LiDAR keyframes [8], generate sparse pose graphs [5, 6], and select navigation anchor points [22]. Streaming submodular algorithms [23, 24] have been used to summarize large image datasets, but have not been used for LiDAR map summarization or SLAM to the best of our knowledge.
@@ III Preliminaries
The marginal value of a set function given an added element is defined as $\Delta f(A\cup\{i\})=f(A\cup\{i\})-f(A)$ where the element $i\in E$ is added to set $A$.
@@
A set function $f:{S}\rightarrow\mathbb{R}$ is submodular if for $A\subseteq B\subseteq E$ and any element $i\in E\,\backslash\,B$ then
$f(A\cup\{i\})-f(A)\geq f(B\cup\{i\})-f(B).$ (2)
Put simply, submodularity is equivalent to diminishing returns.
@@
We conclude with a discussion on streaming submodular algorithms which have proven to be effective at solving combintorial optimization problems with a single pass over the data rather than multiple greedy passes. The streaming algorithm most pertinent to this work is described in Section V-C and given by Algorithm 3, but the intuition is that elements with high marginal value are added to a potential solution based on a guess of the optimal function value. With a range of optimal value guesses, the streaming algorithm can build several potential solution sets and return the best which is guaranteed to be within a 1/2-$\epsilon$ factor of the optimal set where $\epsilon$ is a granularity constant [23].
Notation. We call the set of all collected point clouds E. We denote the set of keyframes as $\mathcal{K}\subseteq\textbf{E}$, and the local submap as $\mathcal{S}\subseteq\mathcal{K}$. All sets have a defined ordering function where E and $\mathcal{K}$ are ordered chronologically and the ordering of $\mathcal{S}$ is defined by when an element is added. We use a matching upper case letter to indicate a member of a set, e.g., $S\in\mathcal{S}$. The position of an element is given by a subscript index $S_{i}$, and we adopt the notation $\mathcal{S}_{<i}$ to denote the subset of $\mathcal{S}$ such that only elements with index less than $i$ are included, e.g., $\mathcal{S}=\{S_{0},S_{1},S_{2}\},~{}\mathcal{S}_{<2}=\{S_{0},S_{1}\}$.
@@ IV LiDAR Scan Neural Network
We adopt the neural network proposed by [7]; the overall framework is presented in Fig. 2. Compared to point- and discretized-based methods [27, 28, 29], the neural network is compact and generalizable because it can achieve fast execution without any semantic information. Instead of using the overlap between the projected range images to supervise the training, we measure the similarity between two point clouds $E_{i}$ and $E_{j}$ directly using the 3D Jaccard index [11]. Defined as $J(E_{i},E_{j})={|E_{i}\bigcap E_{j}|}/{|E_{i}\bigcup E_{j}|}$, the Jaccard index is the equivalent of overlap in the 3D point cloud domain to provide a direct measure of point cloud similarity which enhances training results. Unless otherwise specified, the dimension of the output descriptor is $\text{dim}(\mathcal{G})={256}$.
@@ V-A Keyframe Selection via Feature Descriptor
Our approach relies on the neural network discussed in Section IV to efficiently determine the similarity between any two point clouds. We define the marginal value of any point cloud as the minimum Euclidean distance between the descriptor of a point cloud $E$ and any previously selected keyframe $\Delta f(\mathcal{K}\cup E)=\min_{K\in\mathcal{K}}||\phi(E)-\phi(K)||$. This marginal value inherently defines an objective function
$f(\mathcal{K})=\sum_{K_{i}\in\mathcal{K}}\min_{K_{j}\in\mathcal{K}_{<i}}||\phi(K_{i})-\phi(K_{j})||.$ (6)
Although we do not directly maximize Equation 6, we still obtain a suboptimality bound by only selecting keyframes with high marginal value because $f$ is a non-decreasing monotonic submodular set function, as proven in the following proposition.
@@
Algorithm 1 describes our keyframe selection strategy, where a keyframe is selected according to a feature distance condition with threshold $\alpha$, and a degeneracy condition with threshold $\beta$. In addition to a boolean variable representing whether to save a keyframe or not, we also return the added suboptimality (with respect to Equation 6) $\gamma$ such that the marginal value of the new keyframe is $\alpha-\gamma$. In line 1 we find the smallest distance to any keyframe in descriptor space. If the closest distance is greater than $\alpha$, we return true and report no additional suboptimality on line 3. On lines 5-8, we compute degeneracy and save a keyframe if it exceeds the threshold $\beta$. If neither condition is met, Algorithm 1 returns false. By knowing the marginal value of each keyframe at the time of selection, we also maintain a known suboptimality bound which is given in the following proposition.
@@ V-B Submap Generation via Submodular Maximization
$\sum_{c\in\mathcal{C}}\mathcal{E}(\textbf{t},\textbf{q},E,T),$ (7)
@@
Assumption 1 allows us to approximate $\mathcal{H}_{tt}$ as a sum of sub-Hessians. We denote the sub-Hessian associated with keyframe $K_{i}$ as $\partial\mathcal{H}_{tt}[i]$ and approximate it as $\partial\mathcal{H}_{tt}[i]=\mathcal{H}_{tt}(E-\mathcal{S}_{<i},K_{i})$, where $\mathcal{S}_{<i}$ is the submap including all keyframes with index less than $i$, and the subtraction of point clouds indicates removing the intersection of points ($A-B=A-(A\cap B))$. Because each Hessian is the outer product of the Jacobian with itself, the Hessian must be positive semi-definite $\mathcal{H}_{tt}=J^{T}_{t}J_{t}\succeq 0$, which intuitively means each keyframe cannot make Equation 7 less constrained. We define any keyframe which introduces enough new correspondences such that $\partial\mathcal{H}_{tt}[i]\succ 0$ as a constructive keyframe.
@@
$\Delta f(\mathcal{S}\cup K_{i})=\lambda_{min}(\mathcal{H}_{tt}(E,\mathcal{S}\cup K_{i}))-\lambda_{min}(\mathcal{H}_{tt}(E,\mathcal{S}))$
$=\lambda_{min}(\mathcal{H}_{tt}(E,\mathcal{S})+\partial\mathcal{H}_{tt}[i])-\lambda_{min}(\mathcal{H}_{tt}(E,\mathcal{S}))$
$=\min_{||\textbf{u}||=1}\textbf{u}^{T}[\mathcal{H}_{tt}(E,\mathcal{S})+\partial\mathcal{H}_{tt}[i]]\textbf{u}-\min_{||\textbf{v}||=1}\textbf{v}^{T}\mathcal{H}_{tt}(E,\mathcal{S})\textbf{v}$
$\geq\bar{\textbf{u}}^{T}[\mathcal{H}_{tt}(E,\mathcal{S})+\partial\mathcal{H}_{tt}[i]]\bar{\textbf{u}}-\bar{\textbf{u}}^{T}\mathcal{H}_{tt}(E,\mathcal{S})\bar{\textbf{u}}$
$=\bar{\textbf{u}}^{T}\partial\mathcal{H}_{tt}[i]\bar{\textbf{u}}>0,$
where $\bar{\textbf{u}}$ is the minimizer of $\min_{||\textbf{u}||=1}\textbf{u}^{T}[\mathcal{H}_{tt}(E,\mathcal{S})+\partial\mathcal{H}_{tt}[i]]\textbf{u}$, and the final line is true because the keyframe is constructive. The numerator of Definition 4 is the minimum marginal gain attainable, so if the function is monotone increasing, the submodularity ratio must be bounded away from 00 proving Equation 8 is approximately submodular. ∎
Algorithm 2 shows our approach to submap generation, which begins by finding and saving maximum Hessian eigenvalues (lines 3-6). Note that we do not consider keyframes at least twice the sensor range from the current pose as they share no overlapping points. Each greedy iteration (lines 9-23) computes the highest marginal value keyframe, then adds it to the submap and removes its points from the scan such that the following iterations will compute appropriate sub-Hessians. The resulting submap enjoys a suboptimality bound from Definition 5 if it reaches the cardinality constraint or achieves the maximum attainable value of Equation 8 if the marginal value approaches zero. We only approximate Hessians using a percent ($\sim 25\%$) of randomized scan points as previous work has shown this has a minimal effect on the relative magnitude of ICP Hessian eigenvalues [21].
@@
where $A$ and $B$ are Hermitian matrices with ordered eigenvalues such that $\lambda_{1}\geq...\geq\lambda_{n}$. We use Weyl’s $1st$ inequality ($i=n,j=1$) to upper bound the minimum eigenvalue of the Hessian $\lambda_{min}(\mathcal{H}_{tt}+\partial\mathcal{H}_{tt}[i])\leq\lambda_{min}(\mathcal{H}_{tt})+\lambda_{max}(\partial\mathcal{H}_{tt}[i])$. Next, we note that the maximum eigenvalues of the Hessians obtained on line 5 upper bound the sub-Hessian eigenvalues $\lambda_{max}(\mathcal{H}_{tt}(E,K))\geq\lambda_{max}(\partial\mathcal{H}_{tt}[i])$. This is because the maximum eigenvalue is monotone non-decreasing (Weyl’s $2nd$ inequality $i=1,j=n$) and the Hessian obtained on line 5 is the sum of the sub-Hessian with its counterpart formed by points shared by the scan, keyframe, and previous submap. This combined with the minimum eigenvalue upper bound imply an upper bound on the marginal value of adding keyframe $K_{i}$ of $\Delta\lambda_{min}(\mathcal{S}\cup K)\leq\,\lambda_{max}(\mathcal{H}_{tt}(E,K))$. This means any keyframe with an upper bound smaller than the best marginal gain in the current iteration can be ignored.